OpenStudy (anonymous):

Could you please help me with these problems? [9.01] Determine whether the graph of y = −3x2 + 2x − 8 opens up or down and whether it has a maximum or minimum point. Up; Minimum Up; Maximum Down; Minimum Down; Maximum [9.01] Determine whether the graph of y = x2 + 4x − 7 has a maximum or minimum point, then find the maximum or minimum value. Minimum; (-11, -2) Maximum; (-11, -2) Maximum; (-2, -11) Minimum; (-2, -11) [9.02] What are the x-intercept(s) of the graph of y + 20 = x2 − x? (−5, 0) and (4, 0) (5, 0) and (−4, 0) (−5, 0) and (−4, 0) (5, 0) and (4

OpenStudy (anonymous):

\[y = −3x^2 + 2x − 8 \] the leading coefficient is negative (it is \(-3\)) so it opens down

OpenStudy (anonymous):

since it opens down it has a max, but no minimum

OpenStudy (anonymous):

\[y = x^2 + 4x − 7\] positive leading coefficient, it is 1, so it opens up therefore it has a min, not max minimum is at the second coordinate of the vertex first coordinate of the vertex is always \(-\frac{b}{2a}\) which in your case is \(-\frac{4}{2}=-2\) second coordinate is what you get when you replace \(x\) by \(-2\)

OpenStudy (anonymous):

Thanks so much. Now I just have to try to figure out the other one. :/